BOB'S FATHER AGE IS 3 TIMES OLD AS BOB.4 YEARS AGO,HE WAS 4 TIMES OLDER.HOW OLD IS BOB?: BOB'S FATHER AGE IS 3 TIMES OLD AS BOB.4 YEARS AGO,HE WAS 4 TIMES OLDER.HOW OLD IS BOB?
Answer by Fombitz(1174)

B - Bob's age
F- Bob's father's age
1.F=3B
Four years ago, Bob was B-4 years old and his father was F-4.
2.F-4=4(B-4)
Now you have two equations with two unknowns, you can solve it.
Substitute eq. 1 into eq. 2 and solve for B.
2.F-4=4(B-4)
3B-4=4(B-4)
3B-4=4B-16
-B=-12
B=12
Now use equation 1 to solve for F.
1.F=3B
F=3(12)
F=36
Bob is 12 and his father is 36.
Four years ago, Bob was 8 and his father was 32.